Find the work done on a 48,500 g object to move it on a 22.55 m up a 25.50 degrees incline.
1
Expert's answer
2020-11-06T10:13:32-0500


The net work done against gravity only acounts change in the height above the ground. If we take the initial height of the object "h_i = 0" then its final height is (see the figure):


"h_f = 22.55m\times \sin25.5\degree \approx 8.6295m"

Then, the work done is:


"A = mg(h_f - h_i) = mgh_f"

where "m = 48.5g = 0.0485\space kg" is the mass of the object and "g = 9.81m\/s^2" is the gravitational acceleration. Thus, obtain:


"A = 0.0485\times 9.81\times 8.6295 \approx 4.106\space J"

Answer. 4.106 J.